Please explain in detail

Part 2 of the Problem 1 is:

Continue with the problem. Given Pc = 1/2, find the new consumer surplus, producer surplus, and the deadweight loss.= Problem 1. [20 pts) Given D(x) (x – 3)2 is the price, in dollars per unit, that consumers will pay for x units of an item,
Answer #1


기도에 D%) = 11-3)? Sx) For Point 영 equilibrium P() - S(x) (13) 기 | 까g 262 = X 224 Z CM 3 - 8거 consumer surplus 혜 is the equilit Producer Surpluss (p.5) (no dm van sco a)] = $(x-x22n-x) dx P.s n +20) da t n m?). 3 [ ti 17 0 P.s= -1.34